OpenStudy (ar43r):

4^x=16 how to find x?

OpenStudy (id21):

\[4^x=4^2\] x=2

OpenStudy (amistre64):

there are a few ways; but I think they all involve seeing that 16 = 4^2

OpenStudy (amistre64):

4^x = 16 ; ln the sides ln(4^x) = ln(16) ; simplify x ln(4) = ln(16) ; /ln(4) x = ln(16)/ln(4) a calculator will be able to solve this format

OpenStudy (anonymous):

Using a log to base 4 would probably be simpler than log to base e in this case though.